# Centre for Cyber-Physical Systems IISc, Bengaluru, India

### v1

@@ 3.3 Training Procedure of the Neural $\delta$-ISS-CLF and Corresponding Neural Controller
$\begin{bmatrix}A\\ B\end{bmatrix}^{\top}\begin{bmatrix}2\alpha\beta\Lambda&-(\alpha+\beta)\Lambda\\ -(\alpha+\beta)\Lambda&2\Lambda\end{bmatrix}\begin{bmatrix}A\\ B\end{bmatrix}+\begin{bmatrix}\mathcal{L}_{N}^{2}\textbf{I}&0&0&0\\ 0&0&0&0\\ 0&0&0&-\theta_{p}^{\top}\\ 0&0&-\theta_{p}&\textbf{I}\end{bmatrix}\geq 0,$ (13)
holds, where
$A=\begin{bmatrix}\theta_{0}&\ldots&0&0\\ \vdots&\ddots&\vdots&\vdots\\ 0&\ldots&\theta_{p-1}&0\end{bmatrix},B=\begin{bmatrix}0&\textbf{I}\end{bmatrix},$
where $\theta_{0},\ldots,\theta_{p}$ are the weights of the neural network, $\Lambda\in\mathcal{D}_{\geq 0}^{n_{i}},i\in\{1,\ldots,p\}$, where $n_{i}$ denotes number of neurons in $i$-th layer, and $\alpha,$ and $\beta$ are the minimum and maximum slope of the activation functions, respectively.
@@
Consider a dt-CS $\Xi$ with compact state-space ${\mathbf{X}}$ and unknown dynamics $f$. Given a compact input space ${\mathbf{W}}$, we train two neural networks $V_{\theta,b}$ and $g_{\bar{\theta},\bar{b}}$ representing the $\delta$-ISS control Lyapunov function and the controller for the closed-loop system $\Xi_{g}$, such that the loss functions converge as $L(\psi,\eta)=0,L_{v}(\eta)=0$ and $L_{\mathcal{M}}(\psi,\Lambda,\bar{\Lambda})\leq 0$ over the training data sets $\mathcal{X}$ and ${\mathcal{W}}$ obtained using equations (7) and (8). Then, the closed-loop system under the influence of the controller $g_{\bar{\theta},\bar{b}}$ is guaranteed to be $\delta$-ISS as defined in Definition 2.2.
Proof 3.6
The proof can be staged over satisfying conditions for three loss functions.
  • The first loss $L(\psi,\eta)=0$ implies that the SCP has been solved with the minimum value being $\eta_{S}^{*}$. Hence, the controller is trained to ensure that the closed-loop system is $\delta$-ISS with initial conditions belong to $\mathcal{X}$ (that is, finite over data).
  • The second loss $L_{v}(\eta)=0$ implies the satisfaction of Theorem 3.2, which means that the solution of SCP is also valid for ROP. Hence, the closed-loop system is $\delta$ISS for any initial state in ${\mathbf{X}}$ and any external input sequence in ${\mathsf{w}}$.
  • The third loss $L_{\mathcal{M}}(\psi,\Lambda,\bar{\Lambda})\leq 0$ implies that the matrices $\mathcal{M}(\theta,\Lambda)$ and $\mathcal{M}(\bar{\theta},\bar{\Lambda})$ are positive definite, satisfying the Lipschitz continuity assumption of the neural networks with the assumed Lipschitz bounds.
@@
Hence, the satisfaction of the above theorem leads to ensuring that the closed-loop system is $\delta$-ISS under the action of the controller. This completes the proof.
@@
The training process of the neural $\delta$-ISS-CLF and the corresponding controller is described below:
  1. 1.
    Require: Set of sampled data points $\mathcal{X}$ and ${\mathcal{W}}$ as per Equation (8), and the black box model of the system representing $f(x,u)$ following Assumption 1.
  2. 2.
    Fix all hyper parameters $\varepsilon$, $c=[c_{0},c_{1},c_{2},c_{3},c_{l1},c_{l2}],k=[k_{1},k_{2},k_{3},k_{w},k_{h}],\gamma=[\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{w}],\mathcal{L}_{L},\mathcal{L}_{C},\mathcal{L}_{h}$ a priori. Fix the number of epochs $n_{ep}$ a priori as well.
  3. 3.
    Compute the Lipschitz constants $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{w}$. Estimate the Lipschitz constants $\mathcal{L}_{x}$ and $\mathcal{L}_{u}$ using the reverse Weibull distribution [32]. Get the Lipschitz constant $\mathcal{L}$ as defined in Theorem 3.2.
  4. 4.
    Randomly generate several batches from the dataset. Set the number of batches a priori.
  5. 5.
    In each epoch, pass each batch to the networks and get the outputs $V_{\theta,b}(x,\hat{x})$ and $V_{\theta,b}(f(x,g_{\bar{\theta},\bar{b}}(x,w)),f(\hat{x},g_{\bar{\theta},\bar{b}}(\hat{x},\hat{w})))$. Calculate the loss for each batch using the loss functions mentioned in (11), (14) and (15). The cumulative sum of batch losses gives the epoch loss.
  6. 6.
@@
    Utilizing Adam or stochastic gradient descent optimization techniques with a specified learning rate [33], reduce the loss function and update the trainable parameters $\psi,\eta,\Lambda,\bar{\Lambda}$.
  7. 7.
    Repeat steps 5 and 6 until the loss functions converge according to Theorem 3.5. After successful convergence of the training, the neural network acts as a $\delta$-ISS-CLF $V_{\theta,b}$ under the action of the controller $g_{\bar{\theta},\bar{b}}$, and the closed-loop system is ensured to be incrementally input-to-state stable under the influence of the controller.
@@
Note that if the algorithm does not converge successfully, one cannot judge the incremental input to state stability of the closed-loop system with the specified hyperparameters $\varepsilon,c,k,\gamma,\mathcal{L}_{L},\mathcal{L}_{C}$ and reconsider the choice of initial parameters.
@@
In addition, the initial feasibility of condition (13) is required to satisfy the criterion of loss $L_{\mathcal{P}}$ in (14) according to Theorem 3.5. Choosing small initial weights and biases for neurons can ensure this condition.
Remark 3.9 (Computation of functions $\beta$ and $\gamma$ in (3))
The class $\mathcal{KL}$ function $\beta$ and the class $\mathcal{K}_{\infty}$ function $\gamma$ of Definition 2.2 can be computed using a similar procedure once the trained $\delta$-ISS-CLF is obtained as referred to in [28]. The methodology of obtaining the functions will follow the proof mentioned in Appendix Appendix A. Proof of Theorem 2.8.
@@ 4 Case Studies
The proposed procedure to make a system incremental input-to-state stable using a neural network-based controller is shown through the following case studies of a general nonlinear system, the motion of a one-link manipulator, a jet engine in no-stall mode, and the motion of a rotating rigid spacecraft model. All the case studies were performed using PyTorch in Python 3.10 on a machine with a Windows operating system with Intel Core i7-14700 CPU, 32 GB RAM and NVIDIA GeForce RTX 3080 Ti GPU.
Figure 1: (a)Trajectories starting from different initial conditions under different input signals (for all $k\in{\mathbb{N}}_{0}$, ${\mathsf{w}}(k)=0.2\in{\mathbf{W}}$ for blue curve and $\hat{\mathsf{w}}(k)=0.3\in{\mathbf{W}}$ for red curve), (b) The $\delta$-ISS-CLF plot is greater than zero for all $(x,\hat{x})\in{\mathbf{X}}\times{\mathbf{X}}$, (c) The plot of condition (6c) is always negative corresponding to the same input ${\mathsf{w}}=\hat{{\mathsf{w}}}$.
@@ 4.1 Simple Nonlinear System
$\mathsf{x}(k+1)=\mathsf{x}(k)+\tau(a\sin{(\mathsf{x}(k))}+\tan(\mathsf{u}(k)))$ (16)
where $x(k)$ denotes the state of the system at $k$-th instant. The constant $a=0.1$ represents the rate constant of the system. $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-\frac{\pi}{2},\frac{\pi}{2}]$. Moreover, the input set is given as ${\mathbf{W}}=[-1,1]$. We consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=1.001,\mathcal{L}_{u}=0.01$ using the results in [32].
The goal is to synthesize a controller to enforce the system to be $\delta$-ISS. So, we are to synthesize a valid $\delta$-ISS-CLF $V_{\theta,b}$ under the action of the controller $g_{\bar{\theta},\bar{b}}$. To do this, we first fix the training hyper-parameters as $\epsilon=0.00039,\mathcal{L}_{L}=1,\mathcal{L}_{C}=20,k_{1}=0.00001,k_{2}=0.5,k_{3}=0.0001,k_{w}=0.01$. So, the Lipschitz constant, $\mathcal{L}$, according to Theorem 3.2 is $3.25$. We fix the structure of $V_{\theta,b}$ as $l_{f}=1,h_{f}^{1}=40$ and $g_{\bar{\theta},\bar{b}}$ as $l_{c}=1,h_{c}^{1}=15$.
@@
The successful runs of the algorithm have an average convergence time of 45 minutes.
As can be seen in Figure 1(a), under the different input conditions, the trajectories starting from different initial conditions maintain the same distance after some time instances under the influence of the controller. The $\delta$-ISS-CLF plot for this case is shown in Figure 1(b). Note that, in Figure 1(c), the plot of condition (6c) is shown under the same input condition, and it is always negative, satisfying the condition.
@@ 4.2 One-Link Manipulator
We consider a single link manipulator dynamics [34, 35], whose discrete-time dynamics is governed by the following set of equations:
@@
where $\mathsf{x}_{1},\mathsf{x}_{2}$ denotes angular position and velocity respectively. The constants $M=1,b=0.1$ stand for the mass and damping coefficient of the system, respectively, while $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-\frac{\pi}{4},\frac{\pi}{4}]\times[-\frac{\pi}{4},\frac{\pi}{4}]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-0.5,0.5]$. Also, we consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=1.01,\mathcal{L}_{u}=0.01$.
The goal is to synthesize a controller to enforce the system to be $\delta$-ISS. To do this, we first fix the training hyper-parameters as $\epsilon=0.0157,\mathcal{L}_{L}=1,\mathcal{L}_{C}=40,k_{1}=0.00001,k_{2}=0.5,k_{3}=0.0001,k_{w}=0.01$. So, the Lipschitz constant according to Theorem 3.2 is $3.663$. We fix the structure of $V_{\theta,b}$ as $l_{f}=1,h_{f}^{1}=40$ and $g_{\bar{\theta},\bar{b}}$ as $l_{c}=1,h_{c}^{1}=15$. The training algorithm converges to obtain $\delta$-ISS-CLF $V_{\theta,b}$ along with $\eta=-0.0579$. Hence, $\eta+\mathcal{L}\epsilon=-0.0579+3.663\times 0.0157=-0.00039$, thus by utilizing Theorem 3.5, we can guarantee that the obtained $\delta$-ISS-CLF $V_{\theta,b}$ is valid and that the closed-loop system is assured to be incrementally input-to-state stable under the influence of the controller $g_{\bar{\theta},\bar{b}}$.
@@ 4.3 Jet Engine Model
where $\mathsf{x}_{1}=\mu-1,\mathsf{x}_{2}=\zeta-\rho-2$ with $\mu,\zeta,\rho$ denote the mass flow, the pressure rise and a constant respectively. $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-0.25,0.25]\times[-0.25,0.25]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-0.5,0.5]$. Also, we consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=0.93,\mathcal{L}_{u}=0.01$.
The goal is to synthesize a controller to enforce the system to be $\delta$-ISS. To do this, we first fix the training hyper-parameters as $\epsilon=0.0157,\mathcal{L}_{L}=1,\mathcal{L}_{C}=10,k_{1}=0.00001,k_{2}=0.5,k_{3}=0.001,k_{w}=0.01$. So, the Lipschitz constant $\mathcal{L}$ according to Theorem 3.2 is $3.351$. We fix the structure of $V_{\theta,b}$ as $l_{f}=1,h_{f}^{1}=40$ and $g_{\bar{\theta},\bar{b}}$ as $l_{c}=1,h_{c}^{1}=15$. The training algorithm converges to obtain $\delta$-ISS-CLF $V_{\theta,b}$ along with $\eta=-0.0169$. Hence, $\eta+\mathcal{L}\epsilon=-0.0169+3.351\times 0.005=-0.000145$, therefore, using Theorem 3.5, we can guarantee that the obtained $\delta$-ISS-CLF $V_{\theta,b}$ is valid and that the closed-loop system is assured to be incrementally input-to-state stable under the influence of the controller $g_{\bar{\theta},\bar{b}}$.

### v2

@@ 3.2 Formulation of Loss functions and Training procedure
$\begin{bmatrix}A\\ B\end{bmatrix}^{\top}\!\begin{bmatrix}2\alpha\beta\Lambda&\!\!\!-(\alpha\!+\!\beta)\Lambda\\ -(\alpha\!+\!\beta)\Lambda&2\Lambda\end{bmatrix}\!\begin{bmatrix}A\\ B\end{bmatrix}\!+\!\begin{bmatrix}\mathcal{L}_{N}^{2}\textbf{I}&0&0&0\\ 0&0&0&0\\ 0&0&0&-\theta_{p}^{\top}\\ 0&0&-\theta_{p}&\textbf{I}\end{bmatrix}\!\geq\!0,$ (12)
holds, where $A=\begin{bmatrix}\theta_{0}&\ldots&0&0\\ \vdots&\ddots&\vdots&\vdots\\ 0&\ldots&\theta_{p-1}&0\end{bmatrix},B=\begin{bmatrix}0&\textbf{I}\end{bmatrix}$, $\theta_{0},\ldots,\theta_{p}$ are the weights of the neural network, $\Lambda\in\mathcal{D}_{\geq 0}^{n_{i}},i\in\{1,\ldots,p\}$, where $n_{i}$ denotes number of neurons in $i$-th layer, and $\alpha,$ and $\beta$ are the minimum and maximum slope of the activation functions, respectively.
@@
Consider a dt-CS $\Xi$ with compact state-space $\mathbf{X}$} and unknown dynamics $f$. Given a compact input space $\mathbf{W}$}, we train two neural networks $V_{\theta,b}$ and $g_{\bar{\theta},\bar{b}}$ representing the $\delta$-ISS CLF and the controller for the closed-loop system $\Xi_{g}$, such that the loss functions converge as $L(\psi,\eta)=0,L_{v}(\eta)=0$ and $L_{\mathcal{M}}(\psi,\Lambda,\bar{\Lambda})\leq 0$ over the training data sets $\mathcal{X}$ and $\mathcal{W}$}. Then, the closed-loop system under the influence of the controller $g_{\bar{\theta},\bar{b}}$ is guaranteed to be $\delta$-ISS as defined in Definition 2.2.
Proof:
@@
The first loss $L(\psi,\eta)=0$ implies that the SCP has been solved with optimal $\eta_{S}^{*}$, ensuring the closed-loop system to be $\delta$-ISS with $x\in\mathcal{X}$. The second loss $L_{v}(\eta)=0$ implies the satisfaction of Theorem 3.2, ensuring $\delta$ -ISS for any initial state in $\mathbf{X}$}. The third loss $L_{\mathcal{M}}(\psi,\Lambda,\bar{\Lambda})\leq 0$ implies the neural networks adhere to the Lipschitz bounds, satisfying Assumption 4. Hence, the satisfaction of the above theorem leads to ensuring that the closed-loop system is $\delta$-ISS under the action of the controller. This completes the proof. $\square$
@@
The training process of the neural $\delta$-ISS-CLF and the corresponding controller is described in Algorithm 1.
Algorithm 1 Training of the Neural Networks
1:Black box model of the system $f(x,u)$, Data sets: $\mathcal{X},\mathcal{W}$
[…]
3:Select the hyperparameters $\epsilon$, $c=[c_{0},c_{1},c_{2},c_{3},c_{l1},c_{l2}],k=[k_{1},k_{2},k_{3},k_{w}],\Gamma=[\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{w}],\mathcal{L}_{L},\mathcal{L}_{C},\mathcal{L}_{h}$ and number of epochs.
4:Computation of Lipschitz constants $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{w}$.
5:Estimation of Lipschitz constants $\mathcal{L}_{x}$ and $\mathcal{L}_{u}$ [34]. Compute $\mathcal{L}$ using Theorem 3.2.
6:Initialize Neural networks and trainable parameters $\theta,b,\bar{\theta},\bar{b},\Lambda,\bar{\Lambda},\eta$.
7:for $i\leq Epochs$ (Training starts here) do
8:  Create batches of training data from $\mathcal{X},\mathcal{W}$
9:  Find batch losses using (10), (13) and (14).
@@
10:  Use ADAM or SGD optimizer with specified learning rate [35] to reduce loss and update the trainable parameters.
11:  if Theorem 3.5 is satisfied then
12:   break
13:  end if
14:end for
15:return $V_{\theta,b},g_{\bar{\theta},\bar{b}},\eta$
@@
Note that if the algorithm does not converge successfully, one cannot judge the $\delta$-ISS of the closed-loop system with the specified hyperparameters $c,k,\Gamma,\mathcal{L}_{L},\mathcal{L}_{C}$. To improve convergence, one can incorporate several strategies like reducing the discretization parameter $\epsilon$ [36] or adjusting neural network hyperparameters (architecture, learning rate) [37] or reconsidering the initial choice of hyperparameters of the algorithm as well. Also, the theoretically ideal case of the convergence of the algorithm is that the loss $L(\psi,\eta)$ should be exactly zero; however, in practice, this is generally infeasible. One can consider that the algorithm has converged based on some very small residual margin in the order of $10^{-6}$ to $10^{-4}$ during implementation.
@@
In addition, the initial feasibility of condition (12) is required to satisfy the criterion of loss $L_{\mathcal{M}}$ in (13) according to Theorem 3.5. Choosing small initial weights and biases for neurons can ensure this condition. The choice of Lipschitz bounds of the networks is generally kept small, so that the networks are robust enough corresponding to the inputs.
@@ 4 Case Studies
The proposed procedure to make a system incremental input-to-state stable using a neural network-based controller is shown through the following case studies of a general nonlinear system and the motion of a rotating rigid spacecraft model. All the case studies were performed using PyTorch in Python 3.10 on a machine with a Windows operating system with Intel Core i7-14700 CPU, 32 GB RAM and NVIDIA GeForce RTX 3080 Ti GPU.
Figure 1: (a)Trajectories starting from different initial conditions under same input signals, (b) Trajectories starting from different initial conditions under different input signals (${\mathsf{w}}(k)=0.2$ for red curves, ${\mathsf{w}}(k)=-0.1$ for blue curves) (c) The $\delta$-ISS-CLF plot is greater than zero for all $(x,\hat{x})\in{\mathbf{X}}\times{\mathbf{X}},x\neq\hat{x}$ and equal to zero while $x=\hat{x}$, (d) The Lyapunov value over the trajectories is decaying with time with the heatmap showing Lyapunov level sets.
@@ 4.1 Simple Nonlinear System
$\mathsf{x}(k+1)=\mathsf{x}(k)+\tau(a\sin{(\mathsf{x}(k))}+\tan(\mathsf{u}(k))),$ (15)
where $x(k)$ denotes the state of the system at $k$-th instant. The constant $a=0.1$ represents the rate constant of the system. $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-\frac{\pi}{2},\frac{\pi}{2}]$. Moreover, the input set is given as ${\mathbf{W}}=[-1,1]$. We consider the model to be unknown. However, the Lipschitz constants $\mathcal{L}_{x}=1,\mathcal{L}_{u}=0.01$ are known for the system.
The goal is to synthesize a controller to enforce the system to be $\delta$-ISS. So, we are to synthesize a valid $\delta$-ISS-CLF $V_{\theta,b}$ under the action of the controller $g_{\bar{\theta},\bar{b}}$. To do this, we first fix the training hyper-parameters as $\epsilon=0.00039,\mathcal{L}_{L}=1,\mathcal{L}_{C}=20,{k=[0.00001,0.5,0.0001,0.01],\Gamma=[2,2,2,2]}$. So, the Lipschitz constant, $\mathcal{L}$, according to Theorem 3.2 is $3.25$. We fix the structure of $V_{\theta,b}$ as $l_{v}=1,h_{v}^{1}=40$ and $g_{\bar{\theta},\bar{b}}$ as $l_{c}=1,h_{c}^{1}=15$. The activation function for both $\delta$-ISS-CLF and the controller is considered to be ‘ReLU’.
@@
The successful runs of the algorithm including the dataset generation time have an average convergence time of 45 minutes.
As can be seen in Figure 1(a), under the same input conditions, the trajectories starting from different initial conditions merge towards each other asymptotically under the influence of the controller. The $\delta$-ISS-CLF plot for this case is shown in Figure 1(b). Note that, in Figure 1(c), The Lyapunov level sets are shown using a heatmap plot, while in the same figure, it is shown that the Lyapunov trajectory is decaying over time.
@@ 4.2 One-Link Manipulator
We consider a single link manipulator dynamics [38, 39], whose discrete-time dynamics is given as:
@@
where $\mathsf{x}_{1},\mathsf{x}_{2}$ denotes angular position and velocity respectively. The constants $M=1,b=0.1$ stand for the mass and damping coefficient of the system, respectively, while $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-\frac{\pi}{4},\frac{\pi}{4}]\times[-\frac{\pi}{4},\frac{\pi}{4}]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-0.5,0.5]$. Also, we consider the model to be unknown. However, the Lipschitz constants $\mathcal{L}_{x}=1.01,\mathcal{L}_{u}=0.01$ are known for the system.
The goal is to synthesize a controller to enforce the system to be $\delta$-ISS. To do this, we first fix the training hyper-parameters as $\epsilon=0.0157,\mathcal{L}_{L}=1,\mathcal{L}_{C}=40,k_{1}=0.00001,k_{2}=0.5,k_{3}=0.0001,k_{w}=0.01$. So, the Lipschitz constant according to Theorem 3.2 is $3.663$. We fix the structure of $V_{\theta,b}$ as $l_{v}=1,h_{v}^{1}=40$ and $g_{\bar{\theta},\bar{b}}$ as $l_{c}=1,h_{c}^{1}=15$. The training algorithm converges to obtain $\delta$-ISS-CLF $V_{\theta,b}$ along with $\eta=-0.0579$. Hence, $\eta+\mathcal{L}\epsilon=-0.0579+3.663\times 0.0157=-0.00039$, thus by utilizing Theorem 3.5, we can guarantee that the obtained $\delta$-ISS-CLF $V_{\theta,b}$ is valid and that the closed-loop system is assured to be incrementally input-to-state stable under the influence of the controller $g_{\bar{\theta},\bar{b}}$.
@@ 4.3 Jet Engine Model
where $\mathsf{x}_{1}=\mu-1,\mathsf{x}_{2}=\zeta-\rho-2$ with $\mu,\zeta,\rho$ denote the mass flow, the pressure rise and a constant, respectively. $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-0.25,0.25]\times[-0.25,0.25]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-0.5,0.5]$. Also, we consider the model to be unknown. However, the Lipschitz constants $\mathcal{L}_{x}=0.93,\mathcal{L}_{u}=0.01$are known for the system.
The goal is to synthesize a controller to enforce the system to be $\delta$-ISS. To do this, we first fix the training hyper-parameters as $\epsilon=0.0157,\mathcal{L}_{L}=1,\mathcal{L}_{C}=10,k_{1}=0.00001,k_{2}=0.5,k_{3}=0.001,k_{w}=0.01$. So, the Lipschitz constant $\mathcal{L}$ according to Theorem 3.2 is $3.351$. We fix the structure of $V_{\theta,b}$ as $l_{v}=1,h_{v}^{1}=40$ and $g_{\bar{\theta},\bar{b}}$ as $l_{c}=1,h_{c}^{1}=15$. The training algorithm converges to obtain $\delta$-ISS-CLF $V_{\theta,b}$ along with $\eta=-0.0169$. Hence, $\eta+\mathcal{L}\epsilon=-0.0169+3.351\times 0.005=-0.000145$, therefore, using Theorem 3.5, we can guarantee that the obtained $\delta$-ISS-CLF $V_{\theta,b}$ is valid and that the closed-loop system is assured to be incrementally input-to-state stable under the influence of the controller $g_{\bar{\theta},\bar{b}}$.